OpenStudy (anonymous):

Help in conversion 1800lb/ft into lb/in

OpenStudy (anonymous):

Would I just multiply by 12?

OpenStudy (anonymous):

Or instead I could keep the units in feet and just use .5 if something is 6 inches right?

OpenStudy (dan815):

1 ft = 12 inches so yes!

OpenStudy (dan815):

u must divide

OpenStudy (dan815):

because if you get 1800lb/ft then for 1/12th of the feet how much would you have

OpenStudy (anonymous):

1800*(1/12)=150

OpenStudy (dan815):

ya

OpenStudy (nurali):

1 lb/ft=0.0833 lb/in 1800 lb/ft = 0.0833*1800 lb/in 1800 lb/ft = 150 lb/in

OpenStudy (dan815):

make sense right...

OpenStudy (anonymous):

Ok. But If I wanted to keep the units in feet, could I convert the inches into ft by using (#of inches/12)?

OpenStudy (whpalmer4):

1800lb/ft into lb/in 1800 lb/ft * 1 ft/12 in = 150 lb/in

OpenStudy (whpalmer4):

write your unit conversions as fractions, and make sure the units cancel out properly.

OpenStudy (anonymous):

Ok Thanks everyone